# AutoCrawler : A Progressive Understanding Web Agent for Web Crawler Generation

### v1

@@ Abstract
Web automation is a significant technique that accomplishes complicated web tasks by automating common web actions, enhancing operational efficiency, and reducing the need for manual intervention. Traditional methods, such as wrappers, suffer from limited adaptability and scalability when faced with a new website. On the other hand, generative agents empowered by large language models (LLMs) exhibit poor performance and reusability in open-world scenarios. In this work, we introduce a crawler generation task for vertical information web pages and the paradigm of combining LLMs with crawlers, which helps crawlers handle diverse and changing web environments more efficiently. We propose AutoCrawler , a two-stage framework that leverages the hierarchical structure of HTML for progressive understanding. Through top-down and step-back operations, AutoCrawler can learn from erroneous actions and continuously prune HTML for better action generation. We conduct comprehensive experiments with multiple LLMs and demonstrate the effectiveness of our framework. [^1] Resources of this paper can be found at https://github.com/EZ-hwh/AutoCrawler
@@ 1 Introduction
Web automation refers to the process of programmatically interacting with web-based applications or websites to execute tasks that would typically require human intervention. Web automation streamlines repetitive and time-consuming tasks, significantly enhancing efficiency, accuracy, and scalability across diverse online processes.
In traditional web automation, methods predominantly rely on wrappers, which are scripts or software specifically designed to extract data from predetermined websites or pages. This approach is characteristic of a closed-world scenario, where the automation system only interacts with a pre-defined, limited set of websites or pages and does not extend beyond this specified domain. Consequently, these traditional methods face limitations in adaptability and scalability, struggling to function effectively when encountering new or altered website structures. Given these limitations, both rule-based wrappers and auto wrappers Bronzi et al. (2013), despite their differences, share a common dependency on manually annotated examples for each website Gulhane et al. (2011). For example, more than 1,400 annotations of webpages are used for information extraction Lockard et al. (2019).
Figure 1: Top: HTML have a hierarchical structure DOM tree; Down: Existing web automation framework: Olive arrows refer to handcraft/LLMs prompting process, Violet arrows refer to parser.
The advent of LLMs has revolutionized web automation by introducing advanced capabilities such as planning, reasoning, reflection, and tool use. Leveraging these capabilities, web automation employs LLMs to construct generative agents that can autonomously navigate, interpret, and interact with web content. This effectively solves open-world web-based tasks through sophisticated language understanding and decision-making processes. However, despite these advancements, this paradigm faces two major issues. On one hand, existing web agent frameworks often demonstrate poor performance, with a success rate mentioned as 2.0 Deng et al. (2023) on open-world tasks. On the other hand, a significant weakness encountered in this approach is its insufficient reusability. This implies that these agents are overly dependent on LLMs even when dealing with similar tasks, leading to low efficiency when managing a large volume of repetitive and similar webpages.
@@
In this work, we propose a crawler generation task for vertical information web pages. The goal of this task is to automatically generate a series of predefined rules or action sequences to automatically extract target information. This task calls for a paradigm that LLMs generate crawlers. Compared to traditional wrappers, this paradigm can be quickly adjusted according to different websites and task requirements. This flexibility enables crawlers to handle diverse and changing web environments more efficiently. Compared to the generative agent paradigm, it introduces intermediate rules to enhance reusability and reduce the dependency on LLMs when dealing with similar tasks, thereby improving efficiency when handling a large number of web tasks.
Although LLMs possess strong webpage information extraction ablities, there are still the following challenges in generating crawlers for LLMs: First, LLMs are primarily pre-trained on massive corpora of cleansed, high-quality pure text, lacking exposure to markup languages such as HTML. As a result, LLMs exhibit a limited understanding of the complex structures and semantics inherent in HTML. Second, HTML, as a semi-structured data, contains elements of both structured (tags and attributes) and unstructured (textual content), concurrently encompassing multilayered information nested. It augments the complexity of crawler generation. Third, although LLMs excel in comprehending textual content, they still fall short in understanding the structural information of lengthy documents. This indicates a potential challenge in accurately capturing and utilizing the hierarchical structure inherent in long HTML documents.
Therefore, we introduce AutoCrawler , a two-stage framework designed to address the crawler generation task. An overview of AutoCrawler is presented in Figure 2. Our framework leverages the hierarchical structure of HTML for progressive understanding. Specifically, we propose a heuristic algorithm consisting of top-down and step-back operation based on LLMs. It first tries to refine down to the specific node in the DOM tree containing the target information, and then moves up the DOM tree when execution fails. This process can correct erroneous executions and progressively prune irrelevant parts of the HTML content until it successfully executes.
Our contributions can be summarized as follows:
  • We propose the web crawler generation task and the paradigm of utilizing LLMs for generating crawlers, and we make an analysis on extraction efficiency.
  • We introduce AutoCrawler , a two-phase framework with progressive understanding to generate executable action sequences.
  • Comprehensive experimental results demonstrate the effectiveness of our framework in the web crawler generation task.
@@ 2 Preliminaries
Figure 2: Our framework for the generation of crawler. Left: Progressive generation process, consists of a cycle of top-down and step-back to progressively generate an executable action sequence; Right: Synthesis process, generates a stable action sequence generated based on ones from seed websites.
In this section, we first define the crawler generation task and then present the dataset collection process and its corresponding evaluation metrics.
@@ 2.1 Task Formulation
First, we formulate our crawler generation task. Given a set of webpages on the same website $w\in\mathcal{W}$ describing a subject entity $s$ (also called a topic entity in the previous literature), and its corresponding predefined target attribute $r\in\mathcal{R}$, the task objective is to generate an executable rule/action sequence $\mathcal{A}$ to extract target information $o$ from all webpages.
@@ 2.2 Datasets
We adopt the semi-structure information extraction task as a testbed for the crawler generation task.
@@
Hao et al. (2011) is a Structured Web Data Extraction dataset that contains webpages and golden labels from 80 websites in 8 domains, with 124,291 webpages. Each of the websites from the same domains focuses on 3-5 attributes in the webpages.
@@
We transform the dataset with the following settings. First, we design instructions for each of the domains, and for each of the attributes as the input information for LLMs [^2] Further details about the prompt is in Appendix D.1. Second, for each website in each domain, we sample 100 webpages as the whole test set. We consider the set of webpages on the same websites and the corresponding extraction instruction as a case. For example, for the ESPN websites [^3] https://global.espn.com/nba/ in NBA player domains, the sampled 100 detail webpage of players and the instruction Please extract the team of the player he plays now is a complete test case of our crawler generation task. Third, we pre-process the websites by removing those elements in a webpage that do not contribute to the semantics. We filter out all DOM element nodes with <script> and <style>, as well as delete all attributes in the element node except @class. We replace the original escape characters in the annotations to ensure consistency with the corresponding information on the web.
Ultimately, we collect three datasets containing 320 / 294 / 83 test cases, covering 32 / 221 / 11 different extraction tasks, and comprising a total of 61,566 webpages from 10 different domains.
@@ 2.3 Evaluation Metrics
A single generation from an LLM is capable of directly extracting value from web pages and generating sequences of actions. However, the existing evaluation schemes for web page extraction tasks still follow the traditional metrics of text information extraction tasks, namely precision, recall, and F1 score. They limit the assessment of methods for the crawler generation task to two aspects. First, it focuses on extraction with a single webpage, rather than considering the generalizability from the perspective of a collection of webpages. Second, it does not effectively measure the transferability when adopting the action sequence to other webpages.
To address this issue, we transform the traditional IE task evaluation into an executable evaluation. Based on the traditional IE evaluation on a collection of webpages, we categorize the executability of action sequences into the following six situations. Specifically, for each extraction task on a website, the result is classified based on the extraction result on precision, recall, and f1-score.
(1) Correct, both precision, recall and f1-score equal 1, which indicates the action sequence is precisely; (2) Precision(Prec.), only precision equals 1, which indicates perfect accuracy in the instances extracted following the action sequence, but misses relevant instances; (3) Recall(Reca.), only recall equals 1, which means that it successfully identifies all relevant instances in the webpage but incorrectly identifies some irrelevant instances; (4) Un-executable(Unex.), recall equals 0, which indicates that the action sequence fails to identify relevant instances; (5) Over-estimate(Over.), precision equals 0, which indicates that the action sequence extracts the instances while ground truth is empty; (6) Else: the rest of the situation, including partially extracting the information, etc.

### v2

# AutoScraper: A Progressive Understanding Web Agent for Web Scraper Generation

@@ Abstract
Web scraping is a powerful technique that extracts data from websites, enabling automated data collection, enhancing data analysis capabilities, and minimizing manual data entry efforts. Existing methods, wrappers-based methods suffer from limited adaptability and scalability when faced with a new website, while language agents, empowered by large language models (LLMs), exhibit poor reusability in diverse web environments. In this work, we introduce the paradigm of generating web scrapers with LLMs and propose AutoScraper, a two-stage framework that can handle diverse and changing web environments more efficiently. AutoScraper leverages the hierarchical structure of HTML and similarity across different web pages for generating web scrapers. Besides, we propose a new executability metric for better measuring the performance of web scraper generation tasks. We conduct comprehensive experiments with multiple LLMs and demonstrate the effectiveness of our framework. Our work is now open-source. [^1] Resources of this paper can be found at https://github.com/EZ-hwh/AutoScraper
@@ 1 Introduction
Web scraping is a process where software automates the extraction of data from websites, typically using bots or web scrapers to gather specific information Thapelo et al. (2021). It is important because it allows for efficient data collection and aggregation, which can be crucial for market research, competitive analysis, and real-time data monitoring.
Due to the diversity of sources and information on the internet, the construction of a web scraper requires substantial human effort. Consequently, two types of methods for automatic web information acquisition have been proposed, categorized as wrapper-based and language-agent-based Sarkhel et al. (2023). The wrapper-based method entails complex sequences of operations within customized rule-based functions, which are designed to efficiently access and retrieve desired data from websites, which is especially beneficial for structured websites with stable layouts Kushmerick (1997); Dalvi et al. (2011); Bronzi et al. (2013). Conversely, the language-agent-based method leverages powerful natural language processing capabilities of large language models (LLMs) to interpret free-text queries and directly extract data within websites to meet the demand, effectively handling both structured and dynamic web content Whitehouse et al. (2023); Marco Perini (2024).
Figure 1: An illustration of comparing wrapper-based methods, language-agent-based methods and AutoScraper .
Although both types of methods facilitate web scraping to varying degrees, as shown in Figure 1, they exhibit significant shortcomings in terms of scalability. Wrapper-based method, while reusable, struggles with entirely new website structures, which necessitates extensive human effort to develop additional customized functions Gulhane et al. (2011); Lockard et al. (2019). Conversely, although language-agent-based methods demonstrate superior performance in adapting to new content, their reliance on a limited number of super-powerful API-based LLMs for web scraping incurs considerable time and financial costs. Together, these challenges impede the broader adoption and scalability of current web scraping technologies, limiting their practicality in dynamic and diverse web environments.
To address the shortcomings of the aforementioned two paradigms, the paradigm of generating web scrapers with LLMs would be the optimal solution. On one hand, compared to wrapper-based methods, it fully leverages the reasoning and reflection capacities of LLMs, reducing manual design on new tasks and enhancing scalability. On the other hand, compared to language-agent-based methods, it introduces repeatable extraction procedures, reducing the dependency on LLMs when dealing with similar tasks, and thereby improving efficiency when handling a large number of web tasks. However, there are several challenges associated with using LLMs to generate web scrapers:
  1. 1.
    Long HTML document. Although LLMs excel in comprehending long textual content, HTML, as semi-structured data, comprises both structured (tags and attributes) and unstructured (textual content) elements. Consequently, it is challenging for LLMs to generate executable web scrapers that strictly adhere to the hierarchical structure of web pages in complex markup contexts.
  2. 2.
    Reusability. A good scraper needs to be reusable across multiple web pages. However, the differences in content and structure between various web pages can lead to the creation of a scraper that references a webpage, which can only be applied to some web pages.
  3. 3.
    Appropriate evaluation metrics. For a scraper to be considered useful, it must be able to automatically extract the desired results from all web pages. However, existing evaluation metrics for web information extraction, which focus on the extraction results from individual web pages, do not adequately reflect the usability of the scraper. This can potentially mislead experimental conclusions.
We introduce AutoScraper, a two-stage framework to address the web scraper generation task. Illustrated in Figure 2, AutoScraper comprises two main components: progressive generation and synthesis. The progressive generation stage leverages the hierarchical structure of HTML for progressive understanding to address the long HTML document. Subsequently, the synthesis module integrates multiple scrapers generated on different web pages to produce a cohesive, website-specific scraper that functions universally within that site. Besides, we propose a new evaluation metric for web scraper generation tasks, called the executability metric. Unlike traditional information extraction metrics that measure single web pages, this metric measures multiple web pages within a website, accurately reflecting the reliability and reusability of the scraper.
@@
We evaluate AutoScraper on three available datasets with 8 LLMs. On all three datasets, AutoScraper consistently outperforms all baselines and achieves new state-of-the-art results in zero-shot settings. Also, AutoScraper can surpass supervised learning methods. Moreover, AutoScraper demonstrates superior efficiency on large-scale web information extraction tasks. Compared to traditional wrappers, AutoScraper adjusted more quickly according to different websites and task requirements. This flexibility enables scrappers to handle diverse and changing web environments more efficiently. Compared to the language agent paradigm, it introduces intermediate functions to enhance reusability and reduce the dependency on LLMs when dealing with similar tasks, thereby improving efficiency when handling a large number of web tasks.
2 Related Work
Wrapper-based methods for web scraping utilize the hierarchical structure of the webpage. Method of this category includes rule-based Zheng et al. (2008), learning wrappers (i.e a DOM-specific parser that can extract content) Gulhane et al. (2011); Kushmerick (1997); Dalvi et al. (2011), heuristic algorithm Lockard et al. (2018, 2019) and deep learning neural network Lin et al. (2020); Zhou et al. (2021); Li et al. (2022); Wang et al. (2022). These methods demand substantial human involvement, including creating wrapper annotations, applying heuristic scoring rules (such as visual proximity), crafting features for neural network input, and using prior knowledge for verification. Therefore, it is difficult for wrapper-based methods to automatically scale up when facing web scraping tasks across a large number of different websites.
With the emergence of powerful LLMs OpenAI (2023); Touvron et al. (2023), language agents Sumers et al. (2023) now operate in interactive environments, leveraging LLM-based reasoning, grounding, learning, and decision-making. General language agents, such as Chain-of-Thought Wei et al. (2023), Reflexion Shinn et al. (2023), Self-Refine Madaan et al. (2023), and Self-Debug Chen et al. (2023), capitalize on LLMs’ self-reflection capabilities for iterative planning optimization. However, these agents do not effectively utilize web structural features and fail to simplify the web environment after unsuccessful planning attempts, limiting the optimization of subsequent planning.
Current language agents primarily aim to streamline the web environment Sridhar et al. (2023); Gur et al. (2023); Zheng et al. (2024) and develop strategies for planning and interacting with the web Sodhi et al. (2023); Ma et al. (2023). Nevertheless, these frameworks mainly focus on the concept of open-world web simulation environments Shi et al. (2017); Yao et al. (2023); Deng et al. (2023); Zhou et al. (2023), which encompass a broad spectrum of tasks found in real-life scenarios, such as online shopping, flight booking, and software development. These task scenarios are oriented towards individuals and have significantly different requirements for accuracy and efficiency compared to web scraping.
As a result, current language-agent-based methods cannot effectively exploit the HTML structural similarities across multiple web pages, reducing their dependency on LLMs when performing repetitive operations and leading to inefficiencies.
@@ 3 Preliminaries
Figure 2: AutoScraper framework of two phases: (a) progressive generation and (b) synthesis.
In this section, we first define the scraper generation task and then present the dataset collection process and its corresponding evaluation metrics.
@@ 3.1 Task Formulation
First, we formulate our scraper generation task. Given a set of webpages on the same website $w\in\mathcal{W}$ describing a subject entity $s$ (also called topic entity in the previous literature), and its corresponding predefined target attribute $r\in\mathcal{R}$, the task objective is to generate an executable rule/action sequence $\mathcal{A}$ to extract target information $o$ from all webpages.
@@ 3.2 Datasets
We adopt the semi-structure information extraction task as a testbed for the scraper generation task.
@@
Hao et al. (2011) is a Structured Web Data Extraction dataset that contains webpages from 80 websites in 8 domains, with 124,291 webpages. Each of the websites from the same domains focuses on 3-5 attributes in the web pages.
@@
Dataset $\text{Num}_{\text{Case}}$ $\text{Num}_{\text{Task}}$ $\text{Num}_{\text{Web}}$
Swde 320 32 32,000
Extended Swde 294 221 29,400
Ds1 83 11 186
Table 1: The statistic of web scraping task benchmarks. We report the number of the case ($\text{Num}_{\text{Case}}$), the number of the different extraction task ($\text{Num}_{\text{Task}}$) and the total number of webpages ($\text{Num}_{\text{Web}}$).
We transform the dataset with the following settings. First, we design instructions for each of the domains, and for each of the attributes as the input information for LLMs [^2] Further details about the prompt is in Appendix D.1. Second, for each website in each domain, we sample 100 web pages as the whole test set. We consider the set of webpages on the same websites and the corresponding extraction instruction as a case. For example, for the ESPN websites [^3] https://global.espn.com/nba/ in NBA player domains, the sampled 100-detail webpage of players and the instruction Please extract the team of the player he plays now is a complete case of our scraper generation task. Third, we pre-process the web pages by removing irrelevant elements in a webpage. We use open-source BeautifulSoup library [^4] https://beautifulsoup.readthedocs.io and filter out all DOM element nodes with <script> and <style>, as well as delete all attributes in the element node except @class. We replace the original escape characters in the annotations to ensure consistency with the corresponding information on the web. The statistic of the dataset we transformed is shown in Table 1.
@@ 3.3 Evaluation Metrics
Existing evaluation schemes for web page information extraction tasks still follow the traditional metrics of text information extraction tasks, namely precision, recall, and F1 score. They limit the assessment of methods for the scraper generation task to two aspects. First, it focuses on extraction with a single webpage, rather than considering the generalizability from the perspective of a collection of webpages. Second, it does not effectively measure the transferability when adopting the action sequence to other web pages.
To address this issue, we transform the traditional IE task evaluation into an executable evaluation. Based on the traditional IE evaluation on a collection of web pages, we categorize the executability of action sequences into the following six situations. Specifically, for each extraction task on a website, the result is classified based on the extraction result on precision, recall, and f1-score. (1) Correct: both precision, recall and f1-score equal 1, which indicates the action sequence is precisely; (2) Precision(Prec.): only precision equals 1, which indicates perfect accuracy in the instances extracted following the action sequence, but misses relevant instances; (3) Recall(Reca.): only recall equals 1, which means that it successfully identifies all relevant instances in the webpage but incorrectly identifies some irrelevant instances; (4) Un-executable(Unex.): recall equals 0, which indicates that the action sequence fails to identify relevant instances; (5) Over-estimate(Over.): precision equals 0, which indicates that the action sequence extracts the instances while ground truth is empty; (6) Else: the rest of the situation, including partially extracting the information, etc.